OpenStudy (song_of_the_sole):

find the surface area of the sphere with the given dimension. Leave in terms of pi radius of 60m

OpenStudy (song_of_the_sole):

@MandyWestSide @latinC

OpenStudy (latinc):

4*pi*r^2 = SA

OpenStudy (mandywestside):

45238.93m²

OpenStudy (song_of_the_sole):

i dont think thats right @MandyWestSide

OpenStudy (mandywestside):

Do you have choices for the answers?

OpenStudy (mandywestside):

The solution is 14,400 pi m^2!

OpenStudy (wolf1728):

Surface Area = 4 PI r^2

OpenStudy (song_of_the_sole):

yes A 3600 pi m^2 B 7200 pi M^2 C10800 PI M^2 D 14400 PI M^2

OpenStudy (wolf1728):

Surface Area = 4 PI r^2 If r = 60 then Surface Area = 4*PI * 60^2 Surface Area = PI * 3600 *4 Surface Area = PI * 14,400 Looks like D to me

OpenStudy (song_of_the_sole):

wait i got 3600 wouldnt it be a

OpenStudy (wolf1728):

it is 3,600 times 4

OpenStudy (song_of_the_sole):

oh okay

OpenStudy (wolf1728):

thanks for the medal :-)